OpenStudy (anonymous):

@matt101 Last Two Electricity Problems

OpenStudy (matt101):

How do you think you could solve #4?

OpenStudy (matt101):

You got it!

OpenStudy (anonymous):

How would I rearrange the equation, so I am finding the magnitude?

OpenStudy (matt101):

Yes, you are finding the magnitude of one of the charges

OpenStudy (matt101):

\[65={(9 \times 10^9) \times (6 \times 10^{-6}) \times q_2 \over 0.05^2}\] How would you isolate q2?

OpenStudy (matt101):

Yup!

OpenStudy (anonymous):

@matt101 Is it 108.33 repeating?

OpenStudy (matt101):

Check your math - that's an awfully big charge!

OpenStudy (matt101):

That's it!

OpenStudy (matt101):

Yup

OpenStudy (anonymous):

Okay! Thank you so much! How would I do the three parts to number 5?

OpenStudy (matt101):

It's a question of adding forces to come up with the net force. Just use Coulomb's Law as usual!

OpenStudy (anonymous):

Really? Wow, nearly all of these used his law!

OpenStudy (anonymous):

Would the origin be 0 for A?

OpenStudy (anonymous):

@matt101

OpenStudy (matt101):

Yes

OpenStudy (matt101):

Well keep in mind you have 3 charges in this situation. Each particle will experience 2 forces - one as a result of one of the other particles, and one as a result of the OTHER other particle. This means you'll need to add 2 forces to get the net force on each particle. Each force you can find using Coulomb's Law. Just be careful with the directions of the forces!

OpenStudy (anonymous):

Oh gosh. This is confusing already. I get what you're saying but how do I know which numbers to add?